# Random local access for sampling $k$-SAT solutions

### v2

@@ 6 Concluding remarks
With more involved refinements and optimizations of the arguments in this work, the density constraint $d\lesssim 2^{k/400}$ of 1.1 can be substantially improved (to something closer to $d\lesssim 2^{k/50}$). We omit these additional calculations in favor of expositional clarity to highlight our main result: random local access models exist for arbitrary bounded degree $k$-CNFs at exponential clause density. Furthermore, these arguments can also be adapted (in a similar fashion to [13, 19, 6]) to obtain a random local access model for random $k$-CNFs in a comparable density regime.
@@
Appendix A Correctness of Algorithm 6
In this subsection, we show Proposition 4.6 that with high probability, $(\mathsf{Glauber}(\mathsf{pred}_{u}(0)))_{u\in\mathcal{M}}$ faithfully returns the final outcome $(X_{0}(u))_{u\in\mathcal{M}}$ of the systematic scan Glauber dynamics $(X_{t})_{-T\leq t\leq 0}$ initialized at $X_{-T}=Y$. We will later use Theorem 4.3 to show that therefore, when $T$ is set to be sufficiently large, the distribution of $(\mathsf{MarginSample}(u))_{u\in\mathcal{M}}=(\mathsf{Glauber}(\mathsf{pred}_{u}(0)))_{u\in\mathcal{M}}$ is close to the marginal distribution $\nu=\mu_{\mathcal{M}}$.
Proposition A.1.
Fix $t\leq 0$ and let $u=u_{i(t)}$. Suppose $|R|<80dk^{4}\log n$ after the execution of $\mathsf{Glauber}(t)$ (i.e., Line 4 of Algorithm 6 is never triggered). Then $\mathsf{Glauber}(t)$ faithfully returns $X_{t}(u)$, where $(X_{t})_{-T\leq t\leq 0}$ is the systematic scan Glauber dynamics started at $X_{-T}=Y$.
Proof.
The statement clearly holds for $t=-T$ and $u=u_{i(-T)}$. Since $\mathsf{pred}_{u}(-T)=-T$, by Lines 1–2 of Algorithm 6, we have $\mathsf{Glauber}(-T)=Y(u)=X_{-T}(u)$.
Inductively, for $-T<t\leq 0$, assume the proposition for all $-T\leq t^{\prime}<t$. Let $u=u_{i(t)}$. Suppose $|R|<80dk^{4}\log n$ after the execution of $\mathsf{Glauber}(t)$. Observe that, since $\mathsf{pred}_{w}(t)<t$ for all $w\neq u$, $\mathsf{Glauber}(t)$ only makes further calls to $\mathsf{Glauber}(t^{\prime})$ with $t^{\prime}<t$. Thus, by the inductive hypothesis, all further calls of $\mathsf{Glauber}(t^{\prime})$ have correctly returned the outcomes $X_{t^{\prime}}(u_{i(t^{\prime})})$.
We wish to show that the resampled outcome $\mathsf{Glauber}(t)$ follows the marginal distribution $\nu_{u}^{X_{t-1}(\mathcal{M}\setminus\{u\})}$. Per Lines 5–6, we may assume that $\mathsf{Glauber}(t)$ has never been called before, in which case we directly go to Line 7 of Algorithm 6. Lines 7–8 guarantee that with probability $2\theta$, we assign $X_{t}(u)$ to be 0 or 1 uniformly at random. It remains to show that in Lines 9–19, we are able to resample $X_{t}(u)$ from the padding distribution $\nu_{u}^{\mathrm{pad},X_{t-1}(\mathcal{M}\setminus\{u\})}$.
@@ Appendix A Correctness of Algorithm 6
[…]
@@
Recall that we wish to sample from $\nu_{u}^{\mathrm{pad},X_{t-1}(\mathcal{M}\setminus\{u\})}$. Observe that for any partial assignment $\sigma$, $\nu_{u}^{\mathrm{pad},\sigma}$ is a deterministic function of $\nu_{u}^{\sigma}$ (see Definition 4.5). Since $\nu_{u}^{X_{t-1}(\mathcal{M}\setminus\{u\})}=(\nu_{\Psi})_{u}^{\sigma_{\Lambda}}$, we have $\nu_{v}^{\mathrm{pad},X_{t-1}(\mathcal{M}\setminus\{u\})}=(\nu_{\Psi})_{v}^{\mathrm{pad},\sigma_{\Lambda}}$ as well. Thus, it suffices to sample $c\sim(\nu_{\Psi})_{u}^{\mathrm{pad},\sigma_{\Lambda}}=\nu_{u}^{\mathrm{pad},X_{t-1}(\mathcal{M}\setminus\{u\})}$ which was performed in Line 19. This shows that Lines 9–19 draws $X_{t}(u)$ from the padding distribution $\nu_{u}^{\mathrm{pad},X_{t-1}(\mathcal{M}\setminus\{u\})}$, and finishes the proof. ∎
@@
There are not many $2$-trees containing a fixed vertex in a graph of bounded maximum degree. We recall one example upper bound below that will be sufficient for our purposes.
Observation A.4 (see [10, Corollary 5.7]).
Let $G=(V(G),E(G))$ be a graph or hypergraph with maximum degree $D$. Then for every $v\in V$, the number of 2-trees $Z\subseteq V(G)$ containing $v$ with $|Z|=\ell$ is at most $\frac{(eD^{2})^{\ell-1}}{2}$.
@@
Proposition A.5.
Fix $t\leq 0$. Suppose $\theta:=1-\frac{1}{2}\exp\left(\frac{2edk}{2^{\alpha k}}\right)\geq 0.4$ and $236ed^{3}k^{4}\cdot 0.6^{\alpha k}\leq 1/$ 2. Upon the termination of $\mathsf{Glauber}(t)$, for every $\eta\geq 1$, the size of $R$ satisfies
@@
Proof.
Observe that for every $t_{0}\leq t$, $\mathsf{Glauber}(t_{0})$ calls LB-Sample at most $k+1$ times (in Line 7 and Line 12 of Algorithm 6) before possibly going into another subroutine $\mathsf{Glauber}(t_{1})$. Let $A$ denote the set of timestamps $t_{0}$ such that $\mathsf{Glauber}(t_{0})$ was called at least once. Then we have $|R|\leq(k+1)|A|$.
Observe that every $\mathsf{Glauber}(t_{0})$ with $t_{0}<t$ is triggered by some $x\in W$. Moreover, every $x\in W$ triggers at most $k$ subroutines $\mathsf{Glauber}(t_{0})$ in Line 14. Thus, we get that $|A|\leq k|W|+1$, which gives $|R|\leq(k+1)|A|\leq(k^{2}+k)|W|+k+1\leq 2k^{2}|W|+2k$. ∎
Claim A.7.
The maximum degree of $G$ is at most $16k^{2}d-$ 1.
Proof.
@@
$\{\mathsf{pred}_{w}(t_{1}):w\in\operatorname{vbl}(C_{1})\cap\mathcal{M}\}$ =$=\{s_{1},\dots,s_{k_{1}}\}\text{ with }s_{1}<\dots<s_{k_{1}},$
$\{\mathsf{pred}_{w}(t_{2}):w\in\operatorname{vbl}(C_{2})\cap\mathcal{M}\}$ =$=\{s_{1}^{\prime},\dots,s_{k_{2}}^{\prime}\}\text{ with }s_{1}^{\prime}<\dots<s_{k_{2}}^{\prime}.$
@@
Observe that as $t_{2}$ increments from $s_{k_{1}}-2m+1$ to $s_{1}+3m-1$, we have $\{\mathsf{pred}_{w}(t_{2}):w\in\operatorname{vbl}(C_{2})\cap\mathcal{M}\}\neq\{\mathsf{pred}_{w}(t_{2}-1):w\in\operatorname{vbl}(C_{2})\cap\mathcal{M}\}$ only if $\mathsf{pred}_{w}(t_{2})>\mathsf{pred}_{w}(t_{2}-1)$ for some $w\in\operatorname{vbl}(C_{2})\cap\mathcal{M}$. Moreover, since $|\operatorname{vbl}(C_{2})\cap\mathcal{M}|\leq k$ and $|S|\leq 5m$, we know that there are at most $5k$ such numbers $t_{2}$ in $S$, and these numbers have been completely determined by $\operatorname{vbl}(C_{2})$ and $S$ (i.e., by $\phi(C_{1},t_{1})$ and $C_{2}$). These numbers partition $S$ into at most $15k+$ 1 intervals such that $\{\mathsf{pred}_{w}(t_{2}):w\in\operatorname{vbl}(C_{2})\cap\mathcal{M}\}$ is the same for all $t_{2}$ on each interval. Thus, for every fixed $\phi(C_{1},t_{1})$ and $C_{2}$, the set $\{\{\mathsf{pred}_{w}(t_{2}):w\in\operatorname{vbl}(C_{2})\cap\mathcal{M}\}:t_{2}\in S\}$ has size at most $15k+$ 1.
Therefore, given any $\phi(C_{1},t_{1})$, we can pick a neighbor $\phi(C_{2},t_{2})\sim\phi(C_{1},t_{1})$ by first picking $C_{2}$ (which has $\leq kd$ choices) and then picking an element in $\{\{\mathsf{pred}_{w}(t_{2}):w\in\operatorname{vbl}(C_{2})\cap\mathcal{M}\}:t_{2}\in S\}$ (which has $\leq 5k+1$ choices). So the number of possible $\phi(C_{2},t_{2})\sim\phi(C_{1},t_{1})$ in $G$ is at most $kd(5k+1)\leq 6k^{2}d-1$. ∎
Claim A.8.
@@
Proof.
Consider any two clauses $C,C^{\prime}$ such that $u\in\operatorname{vbl}(C)\cap\operatorname{vbl}(C^{\prime})$. Suppose $\phi(C,t),\phi(C^{\prime},t)\in W^{\prime}$. Clearly $\operatorname{vbl}(C)\cap\operatorname{vbl}(C^{\prime})\neq\emptyset$. Moreover, since the timestamps $\mathsf{pred}_{w}(t)$ over all marked variables $w\in\mathcal{M}$ lie in the range $\{t-m+1,\dots,t\}$, we get that the maximum and minimum of $\{\mathsf{pred}_{w}(t):w\in\operatorname{vbl}(C)\cap\mathcal{M}\}\cup\{\mathsf{pred}_{w}(t):w\in\operatorname{vbl}(C^{\prime})\cap\mathcal{M}\}$ differ by at most $m-1<2m$. Thus we have $\phi(C,t)\sim\phi(C^{\prime},t)$ in $G$. ∎
Claim A.9.
Let $W^{\prime}$ be as in A.8. For every $x\in W$, there exists a path $p_{0}\dots p_{\ell}$ in $G[W]$ such that $p_{0}\in W^{\prime}$ and $p_{\ell}=x$.
Proof.
@@
    Suppose first that $t_{1}=t$, so $x=\phi(C,t)$ for some $C$. Let $u=u_{i(t)}$. If $u\in\operatorname{vbl}(C)$, then $x\in W^{\prime}$ and we are done. Inductively, suppose A.9 holds for all $x^{\prime}=\phi(C^{\prime},t)$ that triggers a recursive call earlier than $x$ in Algorithm 6. If $u\notin\operatorname{vbl}(C)$, then by the while loop condition Line 10, there must exist some $x^{\prime}=\phi(C^{\prime},t)\in W$ such that $x^{\prime}$ triggers a recursive call earlier than $x$, and $\operatorname{vbl}(C)\cap\operatorname{vbl}(C^{\prime})\neq\emptyset$. By the inductive hypothesis, there exists a path $p_{0}\dots p_{\ell}$ in $G[W]$ such that $p_{0}\in W^{\prime}$ and $p_{\ell}=x^{\prime}$. Since the maximum and minimum of $\{\mathsf{pred}_{w}(t):w\in\operatorname{vbl}(C)\cap\mathcal{M}\}\cup\{\mathsf{pred}_{w}(t):w\in\operatorname{vbl}(C^{\prime})\cap\mathcal{M}\}$ differ by at most $m-1<2m$, we get that $x^{\prime}\sim x$ in $G$. Therefore we can extend the path $p_{0}\dots p_{\ell}$ with $p_{\ell+1}=x$. This finishes the inductive step.
@@
    Now suppose $x=\phi(C,t_{1})$ with $t_{1}<t$. By the inductive hypothesis, we can assume A.9 for all $t_{0}\in\{t_{1}+1,\dots,t\}$. Let $u_{1}=u_{i(t_{1})}$.
    Suppose first that $u_{1}\in\operatorname{vbl}(C)$. Since $t_{1}<t$, there must exist $t_{0}\in\{t_{1}+1,\dots,t\}$ and $C^{\prime}\in\mathcal{C}$ such that $\phi(C^{\prime},t_{0})$ triggers $\mathsf{Glauber}(t_{1})$, with $u_{1}\in\operatorname{vbl}(C^{\prime})$ and $t_{1}=\mathsf{pred}_{u_{1}}(t_{0})$. Let $y=\phi(C^{\prime},t_{0})$. Clearly $\operatorname{vbl}(C)\cap\operatorname{vbl}(C^{\prime})\neq\emptyset$. Since $t_{0}-m<t_{1}<t_{0}$, we also know that the maximum and minimum of $\{\mathsf{pred}_{w}(t_{1}):w\in\operatorname{vbl}(C)\cap\mathcal{M}\}\cup\{\mathsf{pred}_{w}(t_{0}):w\in\operatorname{vbl}(C^{\prime})\cap\mathcal{M}\}$ differ by at most $2m-1<2m$. Thus we have $x\sim y$ in $G$. By the inductive hypothesis for $t_{0}$, there exists a path $p_{0}\dots p_{\ell}$ in $G[W]$ such that $p_{0}\in W^{\prime}$ and $p_{\ell}=y$. Since $x\sim y$ in $G$, we can extend the path by $p_{\ell+1}=x$.
    Inductively, suppose $u_{1}\notin\operatorname{vbl}(C_{1})$, and A.9 holds for all $t_{0}\in\{t_{1}+1,\dots,t\}$ and for all $x^{\prime}=(C^{\prime},t_{1})$ that triggers a recursive call earlier than $x$. Then there must exist some $x^{\prime}=\phi(C^{\prime},t_{1})\in W$ such that $x^{\prime}$ triggers a recursive call earlier than $x$, and $\operatorname{vbl}(C)\cap\operatorname{vbl}(C^{\prime})\neq\emptyset$. By the inductive hypothesis, there exists a path $p_{0}\dots p_{\ell}\in G[W]$ such that $p_{0}\in W^{\prime}$ and $p_{\ell}=x^{\prime}$. Since the maximum and minimum of $\{\mathsf{pred}_{w}(t_{1}):w\in\operatorname{vbl}(C)\cap\mathcal{M}\}\cup\{\mathsf{pred}_{w}(t_{1}):w\in\operatorname{vbl}(C^{\prime})\cap\mathcal{M}\}$ differ by at most $m-1<2m$, we get that $x^{\prime}\sim x$ in $G$. Thus we can extend the path by $p_{\ell+1}=x$. This finishes the inductive step.
Claim A.10.
@@
Proof.
[…]
With these claims, we can now prove the proposition. Fix $t\leq 0$ and $\eta\geq 1$. Assume $|R|\geq 24dk^{4}(\eta+1)$, which by A.6 implies that $|W|\geq 6dk^{2}(\eta+1)$. By A.9, we know that $W\cap W^{\prime}\neq\emptyset$; by A.7, $G$ has maximum degree $\leq 6dk^{2}-1$. Thus, by a greedy selection, we can find a 2-tree $Z\subseteq W$ containing some element in $W^{\prime}$ such that $|Z|=\eta+1$.
[…]
@@
Since $Z$ is a 2-tree, we know that for every two $x=\phi(C_{1},t_{1}),y=\phi(C_{2},t_{2})\in Z$, we have $\{\mathsf{pred}_{w}(t_{1}):w\in\operatorname{vbl}(C_{1})\cap\mathcal{M}\}\cap\{\mathsf{pred}_{w}(t_{2}):w\in\operatorname{vbl}(C_{2})\cap\mathcal{M}\}=\emptyset$ (as otherwise $C_{1}$ and $C_{2}$ would share a variable, and the union of these two sets would span an interval of size at most $12(m-1)+1=2m-$ 1, meaning that $x\sim y$ in $G$, which contradicts the fact that $Z$ is an independent set in $G$). In particular, the sets

### v3

@@ 6 Concluding remarks
With more involved refinements and optimizations of the arguments in this work, the density constraint $d\lesssim 2^{k/400}$ of Section˜1 can be substantially improved (to something closer to $d\lesssim 2^{k/50}$). We omit these additional calculations in favor of expositional clarity to highlight our main result: random local access models exist for arbitrary bounded degree $k$-CNFs at exponential clause density. Furthermore, these arguments can also be adapted (in a similar fashion to [13, 19, 7]) to obtain a random local access model for random $k$-CNFs in a comparable density regime.
@@
Appendix A Projected Glauber dynamics for sampling $k$-SAT solutions
In this section, we recall some basics of Markov chains and the projected Glauber dynamics Markov chain for sampling $k$-SAT solutions, introduced in [10].
Definition 2.
Algorithm 6 Projected Glauber dynamics
   Output: a random assignment $X_{*}\in\{0,1\}^{\mathcal{M}}$
1:Sample $X_{0}(u)\overset{\textbf{R}}{\leftarrow}\{0,1\}$ uniformly and independently for each $u\in\mathcal{M}$
2:for $t=1,\dots,T$ do
3:  Sample $u\overset{\textbf{R}}{\leftarrow}\mathcal{M}$
[…]
6:Return $X_{*}=X_{T}$
We contrast Algorithm˜6 with the systematic scan version of the projected Glauber dynamics (Algorithm˜2). Instead of resampling a random variable at time $t$, in the systematic scan Glauber dynamics, we always resample the variable $u_{i(t)}$ at time $t$, where $i(t)$ is a deterministic function of $t$.
Appendix B Correctness of Algorithm˜5
In this section, we show Proposition˜4.4 that with high probability, $(\mathsf{Glauber}(\mathsf{pred}_{u}(0)))_{u\in\mathcal{M}}$ faithfully returns the final outcome $(X_{0}(u))_{u\in\mathcal{M}}$ of the systematic scan Glauber dynamics $(X_{t})_{-T\leq t\leq 0}$ initialized at $X_{-T}=Y$. We will later use Theorem˜4.1 to show that therefore, when $T$ is set to be sufficiently large, the distribution of $(\mathsf{MarginSample}(u))_{u\in\mathcal{M}}=(\mathsf{Glauber}(\mathsf{pred}_{u}(0)))_{u\in\mathcal{M}}$ is close to the marginal distribution $\nu=\mu_{\mathcal{M}}$.
Proposition B.1.
Fix $t\leq 0$ and let $u=u_{i(t)}$. Suppose $|R|<80dk^{4}\log n$ after the execution of $\mathsf{Glauber}(t)$ (i.e., Line 4 of Algorithm˜5 is never triggered). Then $\mathsf{Glauber}(t)$ faithfully returns $X_{t}(u)$, where $(X_{t})_{-T\leq t\leq 0}$ is the systematic scan Glauber dynamics started at $X_{-T}=Y$.
Proof B.2.
The statement clearly holds for t=Tt=-T and $u=u_{i(-T)}$. Since $\mathsf{pred}_{u}(-T)=-T$, by Lines 1–2 of Algorithm˜5, we have $\mathsf{Glauber}(-T)=Y(u)=X_{-T}(u)$.
Inductively, for $T<t\leq 0$, assume the proposition for all $T\leq t^{\prime}<t$. Let $u=u_{i(t)}$. Suppose $|R|<80dk^{4}\log n$ after the execution of $\mathsf{Glauber}(t)$. Observe that, since $\mathsf{pred}_{w}(t)<t$ for all $w\neq u$, $\mathsf{Glauber}(t)$ only makes further calls to $\mathsf{Glauber}(t^{\prime})$ with $t^{\prime}<t$. Thus, by the inductive hypothesis, all further calls of $\mathsf{Glauber}(t^{\prime})$ have correctly returned the outcomes $X_{t^{\prime}}(u_{i(t^{\prime})})$.
We wish to show that the resampled outcome $\mathsf{Glauber}(t)$ follows the marginal distribution $\nu_{u}^{X_{t-1}(\mathcal{M}\setminus\{u\})}$. Per Lines 5–6, we may assume that $\mathsf{Glauber}(t)$ has never been called before, in which case we directly go to Line 7 of Algorithm˜5. Lines 7–8 guarantee that with probability $2\theta$, we assign $X_{t}(u)$ to be 0 or 1 uniformly at random. It remains to show that in Lines 9–19, we are able to resample $X_{t}(u)$ from the padding distribution $\nu_{u}^{\mathrm{pad},X_{t-1}(\mathcal{M}\setminus\{u\})}$.
@@ Appendix B Correctness of Algorithm˜5
We now show property (4). For every $w\in V^{\prime}\setminus\{u\}$, we know that $w$ is added to $V^{\prime}$ in Line 15 due to some clause $C$; if $w$ is marked, then by Line 14 and the inductive hypothesis, we know that we have assigned $\sigma_{\Lambda}(w)\leftarrow\mathsf{Glauber}(\mathsf{pred}_{w}(t))=X_{t-1}(w)\in\{0,1\}$. For every $w\in\Lambda\setminus V^{\prime}$, by Line 18, we have assigned $\sigma_{\Lambda}(w)\leftarrow\textsf{LB-Sample}(\mathsf{pred}_{w}(t))$. If $\textsf{LB-Sample}(\mathsf{pred}_{w}(t))\neq\perp$, then we have $\textsf{LB-Sample}(\mathsf{pred}_{w}(t))=X_{\mathsf{pred}_{w}(t)}(w)=X_{t-1}(w)\in\{0,1\}$.
@@
Recall that we wish to sample from $\nu_{u}^{\mathrm{pad},X_{t-1}(\mathcal{M}\setminus\{u\})}$. Observe that for any partial assignment $\sigma$, $\nu_{u}^{\mathrm{pad},\sigma}$ is a deterministic function of $\nu_{u}^{\sigma}$ (see Definition˜4.3). Since $\nu_{u}^{X_{t-1}(\mathcal{M}\setminus\{u\})}=(\nu_{\Psi})_{u}^{\sigma_{\Lambda}}$, we have $\nu_{v}^{\mathrm{pad},X_{t-1}(\mathcal{M}\setminus\{u\})}=(\nu_{\Psi})_{v}^{\mathrm{pad},\sigma_{\Lambda}}$ as well. Thus, it suffices to sample $c\sim(\nu_{\Psi})_{u}^{\mathrm{pad},\sigma_{\Lambda}}=\nu_{u}^{\mathrm{pad},X_{t-1}(\mathcal{M}\setminus\{u\})}$ which was performed in Line 19. This shows that Lines 9–19 draws $X_{t}(u)$ from the padding distribution $\nu_{u}^{\mathrm{pad},X_{t-1}(\mathcal{M}\setminus\{u\})}$, and finishes the proof.
@@
There are not many $2$-trees containing a fixed vertex in a graph of bounded maximum degree. We recall one example upper bound below that will be sufficient for our purposes. {observation}[see [10, Corollary 5.7]] Let $G=(V(G),E(G))$ be a graph or hypergraph with maximum degree $D$. Then for every $v\in V$, the number of 2-trees $Z\subseteq V(G)$ containing $v$ with $|Z|=\ell$ is at most $\frac{(eD^{2})^{\ell-1}}{2}$.
@@
Proposition B.5.
Fix $t\leq 0$. Suppose $\theta:=1-\frac{1}{2}\exp\left(\frac{2edk}{2^{\alpha k}}\right)\geq 0.4$ and $36ed^{3}k^{4}\cdot 0.6^{\alpha k}\leq 1/2$. Upon the termination of $\mathsf{Glauber}(t)$, for every $\eta\geq 1$, the size of $R$ satisfies
@@
Proof B.7.
Observe that for every $t_{0}\leq t$, $\mathsf{Glauber}(t_{0})$ calls LB-Sample at most $k+1$ times (in Line 7 and Line 12 of Algorithm˜5) before possibly going into another subroutine $\mathsf{Glauber}(t_{1})$. Let $A$ denote the set of timestamps $t_{0}$ such that $\mathsf{Glauber}(t_{0})$ was called at least once. Then we have $|R|\leq(k+1)|A|$.
Observe that every $\mathsf{Glauber}(t_{0})$ with $t_{0}<t$ is triggered by some $x\in W$. Moreover, every $x\in W$ triggers at most $k$ subroutines $\mathsf{Glauber}(t_{0})$ in Line 14. Thus, we get that $|A|\leq k|W|+1$, which gives $|R|\leq(k+1)|A|\leq(k^{2}+k)|W|+k+1\leq 2k^{2}|W|+2k$.
Claim 4.
The maximum degree of $G$ is at most $6k^{2}d-1$.
Proof B.8.
@@
[…]
@@
Observe that as $t_{2}$ increments from $s_{k_{1}}-2m+1$ to $s_{1}+3m-1$, we have $\{\mathsf{pred}_{w}(t_{2}):w\in\operatorname{vbl}(C_{2})\cap\mathcal{M}\}\neq\{\mathsf{pred}_{w}(t_{2}-1):w\in\operatorname{vbl}(C_{2})\cap\mathcal{M}\}$ only if $\mathsf{pred}_{w}(t_{2})>\mathsf{pred}_{w}(t_{2}-1)$ for some $w\in\operatorname{vbl}(C_{2})\cap\mathcal{M}$. Moreover, since $|\operatorname{vbl}(C_{2})\cap\mathcal{M}|\leq k$ and $|S|\leq 5m$, we know that there are at most $5k$ such numbers $t_{2}$ in $S$, and these numbers have been completely determined by $\operatorname{vbl}(C_{2})$ and $S$ (i.e., by $\phi(C_{1},t_{1})$ and $C_{2}$). These numbers partition $S$ into at most $5k+1$ intervals such that $\{\mathsf{pred}_{w}(t_{2}):w\in\operatorname{vbl}(C_{2})\cap\mathcal{M}\}$ is the same for all $t_{2}$ on each interval. Thus, for every fixed $\phi(C_{1},t_{1})$ and $C_{2}$, the set $\{\{\mathsf{pred}_{w}(t_{2}):w\in\operatorname{vbl}(C_{2})\cap\mathcal{M}\}:t_{2}\in S\}$ has size at most $5k+1$.
Therefore, given any $\phi(C_{1},t_{1})$, we can pick a neighbor $\phi(C_{2},t_{2})\sim\phi(C_{1},t_{1})$ by first picking $C_{2}$ (which has $\leq kd$ choices) and then picking an element in $\{\{\mathsf{pred}_{w}(t_{2}):w\in\operatorname{vbl}(C_{2})\cap\mathcal{M}\}:t_{2}\in S\}$ (which has $\leq 5k+1$ choices). So the number of possible $\phi(C_{2},t_{2})\sim\phi(C_{1},t_{1})$ in $G$ is at most $kd(5k+1)\leq 6k^{2}d-1$.
Claim 5.
@@
Proof B.9.
Consider any two clauses $C,C^{\prime}$ such that $u\in\operatorname{vbl}(C)\cap\operatorname{vbl}(C^{\prime})$. Suppose $\phi(C,t),\phi(C^{\prime},t)\in W^{\prime}$. Clearly $\operatorname{vbl}(C)\cap\operatorname{vbl}(C^{\prime})\neq\emptyset$. Moreover, since the timestamps $\mathsf{pred}_{w}(t)$ over all marked variables $w\in\mathcal{M}$ lie in the range $\{t-m+1,\dots,t\}$, we get that the maximum and minimum of $\{\mathsf{pred}_{w}(t):w\in\operatorname{vbl}(C)\cap\mathcal{M}\}\cup\{\mathsf{pred}_{w}(t):w\in\operatorname{vbl}(C^{\prime})\cap\mathcal{M}\}$ differ by at most m1<2mm-1<2m. Thus we have $\phi(C,t)\sim\phi(C^{\prime},t)$ in $G$.
Claim 6.
Let $W^{\prime}$ be as in ˜5. For every $x\in W$, there exists a path $p_{0}\dots p_{\ell}$ in $G[W]$ such that $p_{0}\in W^{\prime}$ and $p_{\ell}=x$.
Proof B.10.
@@
    Suppose first that $t_{1}=t$, so $x=\phi(C,t)$ for some $C$. Let $u=u_{i(t)}$. If $u\in\operatorname{vbl}(C)$, then $x\in W^{\prime}$ and we are done. Inductively, suppose ˜6 holds for all $x^{\prime}=\phi(C^{\prime},t)$ that triggers a recursive call earlier than $x$ in Algorithm˜5. If $u\notin\operatorname{vbl}(C)$, then by the while loop condition Line 10, there must exist some $x^{\prime}=\phi(C^{\prime},t)\in W$ such that $x^{\prime}$ triggers a recursive call earlier than $x$, and $\operatorname{vbl}(C)\cap\operatorname{vbl}(C^{\prime})\neq\emptyset$. By the inductive hypothesis, there exists a path $p_{0}\dots p_{\ell}$ in $G[W]$ such that $p_{0}\in W^{\prime}$ and $p_{\ell}=x^{\prime}$. Since the maximum and minimum of $\{\mathsf{pred}_{w}(t):w\in\operatorname{vbl}(C)\cap\mathcal{M}\}\cup\{\mathsf{pred}_{w}(t):w\in\operatorname{vbl}(C^{\prime})\cap\mathcal{M}\}$ differ by at most m1<2mm-1<2m, we get that $x^{\prime}\sim x$ in $G$. Therefore we can extend the path $p_{0}\dots p_{\ell}$ with $p_{\ell+1}=x$. This finishes the inductive step.
@@
    Now suppose $x=\phi(C,t_{1})$ with $t_{1}<t$. By the inductive hypothesis, we can assume ˜6 for all $t_{0}\in\{t_{1}+1,\dots,t\}$. Let $u_{1}=u_{i(t_{1})}$.
    Suppose first that $u_{1}\in\operatorname{vbl}(C)$. Since $t_{1}<t$, there must exist $t_{0}\in\{t_{1}+1,\dots,t\}$ and $C^{\prime}\in\mathcal{C}$ such that $\phi(C^{\prime},t_{0})$ triggers $\mathsf{Glauber}(t_{1})$, with $u_{1}\in\operatorname{vbl}(C^{\prime})$ and $t_{1}=\mathsf{pred}_{u_{1}}(t_{0})$. Let $y=\phi(C^{\prime},t_{0})$. Clearly $\operatorname{vbl}(C)\cap\operatorname{vbl}(C^{\prime})\neq\emptyset$. Since $t_{0}-m<t_{1}<t_{0}$, we also know that the maximum and minimum of $\{\mathsf{pred}_{w}(t_{1}):w\in\operatorname{vbl}(C)\cap\mathcal{M}\}\cup\{\mathsf{pred}_{w}(t_{0}):w\in\operatorname{vbl}(C^{\prime})\cap\mathcal{M}\}$ differ by at most 2m1<2m2m-1<2m. Thus we have $x\sim y$ in $G$. By the inductive hypothesis for $t_{0}$, there exists a path $p_{0}\dots p_{\ell}$ in $G[W]$ such that $p_{0}\in W^{\prime}$ and $p_{\ell}=y$. Since $x\sim y$ in $G$, we can extend the path by $p_{\ell+1}=x$.
    Inductively, suppose $u_{1}\notin\operatorname{vbl}(C_{1})$, and ˜6 holds for all $t_{0}\in\{t_{1}+1,\dots,t\}$ and for all $x^{\prime}=(C^{\prime},t_{1})$ that triggers a recursive call earlier than $x$. Then there must exist some $x^{\prime}=\phi(C^{\prime},t_{1})\in W$ such that $x^{\prime}$ triggers a recursive call earlier than $x$, and $\operatorname{vbl}(C)\cap\operatorname{vbl}(C^{\prime})\neq\emptyset$. By the inductive hypothesis, there exists a path $p_{0}\dots p_{\ell}\in G[W]$ such that $p_{0}\in W^{\prime}$ and $p_{\ell}=x^{\prime}$. Since the maximum and minimum of $\{\mathsf{pred}_{w}(t_{1}):w\in\operatorname{vbl}(C)\cap\mathcal{M}\}\cup\{\mathsf{pred}_{w}(t_{1}):w\in\operatorname{vbl}(C^{\prime})\cap\mathcal{M}\}$ differ by at most m1<2mm-1<2m, we get that $x^{\prime}\sim x$ in $G$. Thus we can extend the path by $p_{\ell+1}=x$. This finishes the inductive step.
Claim 7.
@@
Proof B.11.
This directly follows from Lines 12–14 of Algorithm˜5. Since $\phi(C,t^{\prime})$ triggers a recursion in Line 14, by the condition in Line 12, for all marked variables $w\in\operatorname{vbl}(C)\cap\mathcal{M}$, the function $\textsf{LB-Sample}(\mathsf{pred}_{w}(t^{\prime}))$ does not satisfy $C$.
With these claims, we can now prove the proposition. Fix $t\leq 0$ and $\eta\geq 1$. Assume $|R|\geq 24dk^{4}(\eta+1)$, which by ˜3 implies that $|W|\geq 6dk^{2}(\eta+1)$. By ˜6, we know that $W\cap W^{\prime}\neq\emptyset$; by ˜4, $G$ has maximum degree $\leq 6dk^{2}-1$. Thus, by a greedy selection, we can find a 2-tree $Z\subseteq W$ containing some element in $W^{\prime}$ such that $|Z|=\eta+1$.
Fix any 2-tree $Z\subseteq V(G)$ such that $|Z|=\eta+1$. For every $x=\phi(C,t^{\prime})\in Z$, if $x\in W$, then we know from ˜7 that $\textsf{LB-Sample}(\mathsf{pred}_{w}(t^{\prime}))$ does not satisfy $C$ for all $w\in\operatorname{vbl}(C)\cap\mathcal{M}$. Since the latter happens with probability at most $(1-\theta)^{\alpha k}$, we have
@@
Since $Z$ is a 2-tree, we know that for every two $x=\phi(C_{1},t_{1}),y=\phi(C_{2},t_{2})\in Z$, we have $\{\mathsf{pred}_{w}(t_{1}):w\in\operatorname{vbl}(C_{1})\cap\mathcal{M}\}\cap\{\mathsf{pred}_{w}(t_{2}):w\in\operatorname{vbl}(C_{2})\cap\mathcal{M}\}=\emptyset$ (as otherwise $C_{1}$ and $C_{2}$ would share a variable, and the union of these two sets would span an interval of size at most 2(m1)+1=2m12(m-1)+1=2m-1, meaning that $x\sim y$ in $G$, which contradicts the fact that $Z$ is an independent set in $G$). In particular, the sets